OpenStudy (anonymous):

fifth order taylor polynomial for cos(x) near 0

OpenStudy (anonymous):

so at a=0? that would be a macluarin polynomial...

OpenStudy (athe):

macluarin: \[f(x)=f(0)+\frac{f'(0)}{1!}\cdot x+\frac{f''(0)}{2!}\cdot x^2+\dots+\frac{f'''''(0)}{5!}\cdot x^5, \] the \[f(x)=\cos(x) \]